# DepthART: Monocular Depth Estimation as Autoregressive Refinement Task

### v2

@@ Introduction
Next, we encode ground-truth depth ${\bm{D}}$ with the same VQ-VAE encoder into continuous features $f_{{\bm{D}}}$, discarding quantization process. Residual prediction targets $\{t_{1},t_{2},\dots,t_{K}\}$ can be constructed based on encoded depth features $f_{\bm{D}}$ and a series of models predictions up to current scale. This process is done similarly to VQ-VAE decomposition method (eqs. 2 and 3):
$\delta_{k}=f_{{\bm{D}}}-\sum_{i=1}^{k-1}\eta_{i}(z_{i})$ (6)
@@
As the result we form a new set of training samples, where $z_{k}$ and $t_{k}$ are model inputs and targets respectively. In contrast to original VAR training, these token maps are dynamically constructed at every training step rather than relying on a single predefined VQ decomposition (see fig. 3).
The training process benefits from such formulation in a few ways. Firstly, our procedure enables model self-refinement by making model aware of own predictions and framing the task as residual refinement. Since VQ-VAE (Tian et al. 2024) decomposition comes from approximating continuous features with summation of discrete token maps, multiple plausible decompositions can exist. Secondly, we argue that exploring various ways of decomposing input into token maps is beneficial for model training. The proposed training method eliminates single-mode limitation and facilities multi-modal solutions discovery (see fig. 2).
Depth Autoregressive Transformer
We develop our depth autoregressive transformer by leveraging the visual autoregressive transformer model, which is based on the GPT-2 architecture (Radford et al. 2019) and pretrained through visual autoregressive modeling (Tian et al. 2024) on the ImageNet dataset (Deng et al. 2009). Originally designed for class-conditioned image generation, we repurpose this model for image-conditioned depth map estimation. We emperically found that the optimal performance is achieved by inputting concatenated image tokens with previously predicted depth tokens. The complete training and inference pipeline is illustrated in fig. 3 (right).
@@
GP-2 (EffNet-B5) 0.23 0.175 0.427 0.247 0.162 0.125 0.162 0.121 4.70 12.8 5.2
Midas (ResNeXt-101) 0.203 0.160 0.325 0.207 0.143 0.116 0.140 0.112 2.71 10.8 3.5
AdaBins (EffNet-B5) 0.235 0.184 0.323 0.206 0.141 0.115 0.161 0.125 4.65 12.6 4
DiT-depth (DiT) 0.309 0.220 0.252 0.169 0.149 0.120 0.169 0.127 2.86 8.71 4.3
DPT (ViT-L) 0.198 0.150 0.435 0.251 0.121 0.107 0.134 0.108 2.97 8.14 2.9
VAR (GPT-2) 0.245 0.285 0.396 0.294 0.185 0.141 0.177 0.133 1.98 9.44 6.1
DepthART (Ours, GPT-2) 0.196 0.177 0.275 0.178 0.141 0.115 0.129 0.106 1.91 7.27 2.1
@@
Table 1: Quantitative evaluation across benchmarks not seen during training. Overall performance is summarized using a rank metric. Our depth autoregressive transformer, trained with DepthART, outperforms the original VAR training procedure and achieves the highest overall performance among a diverse set of depth estimation baselines.
Figure 5: Qualitative comparison of point clouds reconstructed from predicted depth maps on the IBIMS dataset. The depth autoregressive transformer trained with DepthART delivers higher-quality reconstructions, particularly in planar regions.
@@
Figure 6: Qualitative comparison of the depth autoregressive transformer trained with DepthART against various baselines. Our model generates more precise depth estimates in planar regions while maintaining the overall scene structure.
Experiments
@@
In this section, we empirically validate the effectiveness of our proposed approach on several depth estimation benchmarks that were not used during training. Specifically, we demonstrate that: (1) the DepthART training method significantly enhances the performance of the depth autoregressive transformer and introduces multimodality compared to the original VAR, (2) the depth autoregressive transformer trained with DepthART achieves comparable or superior accuracy relative to other baselines.
Implementation details
@@
For our experimental evaluation, we chose to predict depth maps up to an unknown scale (SI). Unlike the commonly used scale-and-shift invariant training approach (Ranftl et al. 2020), scale-invariant training preserves the ability to reconstruct geometry from the predicted depth maps, which is essential for practical applications like single-view reconstruction. Therefore, prior to inputting the depth maps into VQ-VAE, we apply the following transformation:
${\bm{D}}_{norm}=\frac{{\bm{D}}}{{\bm{D}}_{98}+\epsilon}\times 2-1$ (9)
where $d_{98}$$98\%$ percentile of individual depth map.
@@
Our depth autoregressive transformer is trained with DepthART using AdamW (Loshchilov and Hutter 2019) optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$ and batch size equals to 4. Additionally we decrease learning rate during training with StepLR scheduler with a step size of $10,000$ and a gamma of $0.8$. Training of our model takes 2 days using 4 NVIDIA H100 GPUs.
Training protocol
@@
To ensure consistent training conditions across all models, we train both the depth autoregressive transformer and baseline models on the same dataset. Due to the requirement of dense ground-truth depth maps for variational autoencoders, we utilize the highly realistic synthetic HyperSim dataset (Roberts et al. 2021), which includes 461 diverse indoor scenes. The pretrained VQ-VAE (Tian et al. 2024) used in our experiments generates multi-scale token maps only up to a maximum resolution of 256$\times$256, so we train all models at this resolution.
Evaluation protocol
@@
Evaluation is performed on four datasets unseen during training: NYUv2 (Silberman et al. 2012) and IBIMS  (Koch et al. 2019) capturing indoor environments, TUM  (Li et al. 2019) capturing dynamic humans in indoor environment, ETH3D (Schops et al. 2017) providing high-quality depth maps for outdoor environments. Since all models trained to predict depth maps up to unknown scale, we first align predictions with ground-truth depth maps in terms of $\mathcal{L}_{1}$. Firstly, we evaluate accuracy of estimated depth maps using two commonly used metrics: Absolute Mean Relative Error (AbsRel) ($\downarrow$) and $\delta_{1}(\downarrow)$. Additionally, we assess the predicted depth maps using depth planar region deviations (pe-fla $\downarrow$) and plane orientation error (pe-ori, in ${}^{\circ},\downarrow$) on IBIMS dataset (Koch et al. 2019).
Baselines
@@
We evaluate our approach against a diverse set of baseline models, organized into three categories. First, we consider several widely used depth estimation architectures trained discriminatively with an $\mathcal{L}_{2}$ regression loss, including MiDaS (Ranftl et al. 2020), GP2 (Patakin et al. 2022), and DPT (Ranftl, Bochkovskiy, and Koltun 2021). We also include AdaBins (Bhat, Alhashim, and Wonka 2021), which represents a classification-based approach to depth estimation. Additionally, we evaluate DiT (Peebles and Xie 2023), a transformer-based diffusion model pretrained on ImageNet. Finally, to assess the impact of our training procedure, we compare our depth autoregressive transformer trained with DepthART against the original VAR approach. More baseline training details are provided in the appendix.
Experimental results
VAR vs DepthART training.
@@
To demonstrate the advantages of our approach, we trained our depth autoregressive transformer on the HyperSim dataset using both the original VAR training procedure and our DepthART method. Figure 4 presents a detailed comparison of these training methods, evaluated on the ETH3D dataset. We assess the reconstruction quality by calculating the AbsRel metric (fig. 4, left) between the intermediate depth maps (fig. 1, bottom) decoded at each autoregression step, based on cumulative predictions from both VAR and DepthART. Since we used a pretrained VQ-VAE without fine-tuning it for depth estimation, we also present its end-to-end depth map reconstruction error as a soft limit achievable by our depth autoregressive transformer. While the model trained using the VAR approach struggles to improve reconstruction quality at early scales, the DepthART-trained model consistently refines its predictions, achieving an overall reduction in relative error of approximately 70% compared to the baseline. Notably, the DepthART-trained model discovers a slightly better decomposition at the second scale than the reference provided by VQ-VAE, highlighting the non-uniqueness and potential suboptimality of the VQ-VAE decomposition. Additionally, we calculate the entropy of the predicted token distributions at each autoregression step (fig. 4, right). The higher entropy, coupled with the improved reconstruction quality, confirms that DepthART facilitates multi-modal training and leads to the discovery of more optimal prediction trajectories.
Comparison against baselines
@@
To prove the efficiency of autoregressive approach in depth estimation, we train a set of popular baselines in similar conditions. Table 1 presents the evaluation results of all models on benchmarks that were not seen during training. To assess overall performance, we calculated each model’s rank on every dataset, and then averaged these ranks across all datasets. As can be seen from the Table 1, the depth autoregressive transformer trained with DepthART achieves the best overall performance. Notably, both models trained with the VAR and DepthART methods showed significantly better planar depth accuracy on the IBIMS dataset. We provide qualitative comparison of predicted depth maps in Figure 6. Besides, point clouds reconstructed from predicted depth maps (fig. 5) further support this observation. These results highlight the potential of autoregressive models for depth estimation.
Discussion
@@
This work demonstrates the potential of generative autoregressive modeling for monocular depth estimation. Currently, our depth autoregressive transformer builds on VAR pretraining, which is constrained by the the ImageNet dataset. We believe that pretraining on more extensive and diverse datasets, such as those used for text-to-image generation, could significantly enhance our model’s performance. A primary limitation of our approach is the reliance on a VQ-VAE network derived from VAR, which has not been fine-tuned or retrained. This VQ-VAE was trained at low resolution on the relatively smaller OpenImages dataset (Kuznetsova et al. 2020), in contrast to larger, more recent datasets like LAION-5B (Schuhmann et al. 2022). We anticipate that upgrading to a higher-quality VQ-VAE could greatly benefit our method, and we identify these limitations as key directions for future research.
Conclusion
@@
In this paper, we tackle the depth estimation problem through an autoregressive lens, specifically adapting the visual autoregressive modeling approach (Tian et al. 2024) for this task. Originally designed for class-conditioned image generation, we repurposed the visual autoregressive transformer for image-conditioned depth map estimation, introducing the Depth Autoregressive Transformer. Our analysis highlights limitations in the standard VAR training process, which leads to suboptimal accuracy on public depth benchmarks. To address these challenges, we proposed a novel training formulation, the Depth Autoregressive Refinement Task (DepthART). The Depth Autoregressive Transformer trained with DepthART showed substantial performance improvements over the VAR procedure and achieved competitive or superior results on public benchmarks compared to recent methods. Our approach enhances the model’s self-refinement ability and resolves the unimodality issues of visual autoregressive modeling, as demonstrated through empirical evaluation.
Appendix
Depth Autoregressive Transformer
Architecture
Originally, authors of VAR trained several models ranging from 300 million of parameters(depth=16) to 1 billion (width depth=24). We choose the smallest model with 300 million of parameters (depth=16) to make our model size compatible with other baselines. Secondly, since we train all models on a single synthethic dataset using lower number of parameters helps to avoid overfitting.
Inference
During the inference stage, we employed the top-k sampling algorithm to process the predicted token distribution. This method involves selecting the top-k token IDs from the distribution and subsequently sampling from this reduced set with replacement, generating outputs based on the multinomial distribution of the selected tokens. The use of top-k sampling allows for a more diverse range of generated tokens, which can help mitigate the deterministic nature of traditional methods such as argmax.
Baseline training
To evaluate the performance of our proposed model, we conducted a series of experiments involving multiple baseline models. All the models listed below were trained from ImageNet pretraining.
DPT, MiDaS, GP2
Three baseline models: the Dense Prediction Transformer (DPT) with a Vision Transformer (ViT) backbone with 344M parameters, the MiDaS model with a ResNeXt-101 backbone with 100M parameters, and the GP2 (EffNet-B5) model with 30M parameters, all optimized using $L_{2}$ loss on the Hypersim dataset at a resolution of $256\times 256$. For these models, we employed the AdamW optimizer, configuring it with a learning rate of $10^{-4}$ and a weight decay of 0.01. We also applied a StepLR scheduler with a step size of 10,000 iterations and a decay factor of 0.8 to manage the learning rate decay effectively.
DiT
A Diffusion Transformer (DiT) model with around 675M parameters, pretrained on ImageNet, was adapted for the task of depth prediction. To leverage the class-conditioned capabilities of the Diffusion Transformer, we concatenated the encoded depth and image latents along the feature dimension before passing them through the first convolutional projection layer. This modification effectively doubled the number of input channels, necessitating the duplication of the input layer’s weight tensor, with the weights halved to maintain the same initialization scale.
AdaBins
We use original model’s architecture with 78M parameters. We adhered to the original training methodology, incorporating the bin-center density loss proposed in the original work alongside the standard pixel-wise depth loss.
VQ-VAE
In this paper we used VQ-VAE derived from VAR paper. We found this model is able to encode depth maps successfully in most case. To encode it we replicate channel dimension to mimic RGB image and normalize depth value to a proper range. Still, We observe an instability issues of VQ-VAE. Specificially, this model can unpredictably corrupt encoding signal that contains errors in some pixels. Despite these values are in the valid range this model fails to reconstruct the underlying depth map from self encoded tokens. We consider these samples as out-of-distribution outliers since the VQ-VAE unlikely saw such differences in adjacent pixels during training. We demonstrate a couple of such outliers below.
Figure 7: By decoding depth maps on each autoregression step (after each sequentially predicted scale), we illustrate the depth refinement process introduced by our DepthART training problem formulation. Despite moderate improvmenets of numerical measures of reconstruction quality on finer scales (see fig.4 left in our paper), they are crucial for revealing precise object boundaries.
Figure 8: Qualitative comparison of point clouds reconstructed from depth predictions on IBIMS dataset
Figure 9: Qualitative comparison of point clouds reconstructed from depth predictions on NYU dataset
Figure 10: Qualitative comparison of depth maps generated by DepthART and baseline models on in-the-wild images from the COCOval2017 dataset. Despite all models were trained on synthetic indoor images and pre-trained on ImageNet, DepthART demonstrates superior generalization by producing more accurate depth predictions for out-of-distribution objects, including animals, cars, and people. Moreover, our model produces smoother depth gradients.

### v3

@@ 4.0.1 Depth Autoregressive Refinement Task
Next, we encode ground-truth depth D with the same VQ-VAE encoder into continuous features $f_{{\textit{D}}}$, discarding quantization process. Residual prediction targets $\{t_{1},t_{2},\dots,t_{K}\}$ can be constructed based on encoded depth features $f_{\textit{D}}$ and a series of models predictions up to current scale. This process is done similarly to VQ-VAE decomposition method (Eq. 2, 3):
$\delta_{k}=f_{{\textit{D}}}-\sum_{i=1}^{k-1}\eta_{i}(z_{i})$ (6)
@@
As the result we form a new set of training samples, where $z_{k}$ and $t_{k}$ are model inputs and targets respectively. In contrast to original VAR training, these token maps are dynamically constructed at every training step rather than relying on a single predefined VQ-VAE decomposition (see Figure 2).
The training process benefits from such formulation in a few ways. Firstly, our procedure enables model self-refinement by making model aware of own predictions and framing the task as residual refinement. Since VQ-VAE Tian et al. (2024) decomposition comes from approximating continuous features with summation of discrete token maps, multiple plausible decompositions can exist. Secondly, we argue that exploring various ways of decomposing input into token maps is beneficial for model training. Thirdly, this training procedure better aligns with the inference process, where the model receives its own outputs at each step, thus reducing the training-inference discrepancy.
@@
GP-2 (EffNet-B5) 13.4G 41.4 / 21.4 30M 0.23 0.175 0.427 0.247 0.162 0.125 0.162 0.121 4.70 12.8 5.2
Midas (ResNeXt-101) 92G 28.1 / 10.9 104.5M 0.203 0.160 0.325 0.207 0.143 0.116 0.140 0.112 2.71 10.8 3.5
AdaBins (EffNet-B5) 317G 26.8 / 13.4 78.2M 0.235 0.184 0.323 0.206 0.141 0.115 0.161 0.125 4.65 12.6 4
DiT-depth (DiT) 3.2T 514.5 / 217.5 675M 0.309 0.220 0.252 0.169 0.149 0.120 0.169 0.127 2.86 8.71 4.3
DPT (ViT-L) 963G 128.7 / 47.6 344M 0.198 0.150 0.435 0.251 0.121 0.107 0.134 0.108 2.97 8.14 2.9
VAR (GPT-2) 1.17T 187.9 / 74.8 310M 0.245 0.285 0.396 0.294 0.185 0.141 0.177 0.133 1.98 9.44 6.1
VAR+DepthART (Ours, GPT-2) 1.17T 192.2 / 74.8 310M 0.196 0.177 0.275 0.178 0.141 0.115 0.129 0.106 1.91 7.27 2.1
@@
Table 1: Quantitative evaluation across benchmarks unseen during training. Overall performance is summarized using a rank metric. VAR trained with DepthART outperforms the original VAR trained with teacher forcing and achieves the highest overall performance among a diverse set of depth estimation baselines. Trainable parameters count, FLOPS, Training iteration time and Inference time on 1×H100 with batch size 1 are also provided.
@@
Figure 4: Qualitative comparison of the VAR trained with DepthART against various baselines. Our model generates more precise depth estimates in planar regions while maintaining the overall scene structure.
Figure 5: Qualitative comparison of point clouds reconstructed from predicted depth maps on the IBIMS dataset. The visual autoregressive transformer trained with DepthART delivers higher-quality reconstructions, particularly in planar regions.
5 Experiments
@@
In this section, we empirically validate the effectiveness of our proposed approach on several depth estimation benchmarks that were not used during training. Specifically, we demonstrate that: (1) the usage of DepthART training method significantly enhances the performance of the Visual Autoregressive Transformer applied to depth estimation task compared to the usage of teacher forcing, (2) the Visual Autoregressive Transformer trained with DepthART achieves comparable or superior accuracy relative to other baselines.
5.0.1 Implementation Details
@@
For our experimental evaluation, we chose to predict depth maps up to an unknown scale (SI). Unlike the commonly used scale-and-shift invariant training approach Ranftl et al. (2020), scale-invariant training preserves the ability to reconstruct geometry from the predicted depth maps, which is essential for practical applications like single-view reconstruction. Therefore, prior to inputting the depth maps into VQ-VAE, we apply the following transformation:
${\textit{D}}_{norm}=\frac{{\textit{D}}}{{\textit{D}}_{98}+\epsilon}\times 2-1$ (9)
where ${\textit{D}}_{98}$$98\%$ percentile of individual depth map.
@@
Visual Autoregressive Transformer is trained with DepthART using AdamW Loshchilov and Hutter (2019) optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$ and batch size equals to 4. Additionally we decrease learning rate during training with StepLR scheduler with a step size of $10,000$ and a gamma of $0.8$. Training of our model takes 17 hours using 4 NVIDIA H100 GPUs.
Training protocol.
@@
To ensure consistent training conditions across all models, we train both the Visual Autoregressive Transformer and baseline models on the same dataset. Due to the requirement of dense ground-truth depth maps for variational autoencoders, we utilize the highly realistic synthetic HyperSim dataset Roberts et al. (2021), which includes 461 diverse indoor scenes. The pretrained VQ-VAE Tian et al. (2024) used in our experiments generates multi-scale token maps only up to a maximum resolution of 256$\times$256, so we train all models at this resolution.
Evaluation protocol.
@@
Evaluation is performed on four datasets unseen during training: NYUv2 Silberman et al. (2012) and IBIMS  Koch et al. (2019) capturing indoor environments, TUM  Li et al. (2019) capturing dynamic humans in indoor environment, ETH3D Schops et al. (2017) providing high-quality depth maps for outdoor environments. Since all models trained to predict depth maps up to unknown scale, we first align predictions with ground-truth depth maps in terms of $\mathcal{L}_{1}$. Firstly, we evaluate accuracy of estimated depth maps using two commonly used metrics: Absolute Mean Relative Error (AbsRel) ($\downarrow$) and $\delta_{1}(\downarrow)$. Additionally, we assess the predicted depth maps using depth planar region deviations (pe-fla $\downarrow$) and plane orientation error (pe-ori, in ${}^{\circ},\downarrow$) on IBIMS dataset Koch et al. (2019). Recent state-of-the-art generative depth estimation methods have shown that large-scale pretraining plays a crucial role in reducing the synthetic-to-real domain gap. For instance, popular generative models like Marigold  Ke et al. (2024) leverage internet-scale pretraining on images before being fine-tuned on small synthetic datasets for depth estimation. This strong image prior significantly decreases the synthetic-to-real domain gap, enabling these generative models to achieve high zero-shot performance on real-world depth estimation benchmarks. Therefore, we do not perform additional steps to address the domain gap.
Baselines.
@@
We evaluate our approach against a diverse set of baseline models, organized into three categories. First, we consider several widely used depth estimation architectures trained discriminatively with an $\mathcal{L}_{2}$ regression loss, including MiDaS Ranftl et al. (2020), GP2 Patakin et al. (2022), and DPT Ranftl et al. (2021). We also include AdaBins Bhat et al. (2021), which represents a classification-based approach to depth estimation. Additionally, we evaluate DiT Peebles and Xie (2023), a transformer-based diffusion model pretrained on ImageNet. Finally, to assess the impact of our training procedure, we compare VAR trained with DepthART (VAR+DepthART) against the VAR trained with original procedure - teacher forcing. More baseline training details are provided in the supplementary materials.
5.1 Experimental Results
5.1.1 Teacher Forcing vs DepthART Training
@@
To demonstrate the advantages of our approach, we trained VAR on the HyperSim dataset using both the original teacher forcing training procedure and our DepthART method. Figure 3 presents a detailed comparison of these training methods, evaluated on the ETH3D dataset. We assess the reconstruction quality by calculating the AbsRel metric (Figure 3) between the intermediate depth maps decoded at each autoregression step, based on cumulative predictions from both teacher forcing and DepthART. Since we used a pretrained VQ-VAE without fine-tuning it for depth estimation, we also present its end-to-end depth map reconstruction error as a soft limit achievable by VAR. While the model trained using the teacher forcing approach struggles to improve reconstruction quality at early scales, the DepthART-trained model consistently refines its predictions, achieving an overall reduction in relative error of approximately 70% compared to the baseline. Notably, the DepthART-trained model discovers a slightly better decomposition at the second scale than the reference provided by VQ-VAE, highlighting the non-uniqueness and potential suboptimality of the VQ-VAE decomposition.
5.1.2 Comparison Against Baselines
@@
To prove the efficiency of autoregressive approach in depth estimation, we train a set of popular baselines in similar conditions. Table 1 presents the evaluation results of all models on benchmarks that were not seen during training. To assess overall performance, we calculated each model’s rank on every dataset, and then averaged these ranks across all datasets. As can be seen from the Table 1, the VAR trained with DepthART achieves the best overall performance. Notably, both VAR models trained with the teacher forcing and DepthART methods showed significantly better planar depth accuracy on the IBIMS dataset. We provide qualitative comparison of predicted depth maps in Figure 4. Besides, point clouds reconstructed from predicted depth maps (Figure 5) further support this observation. These results highlight the potential of autoregressive models for depth estimation.
6 Ablation Studies
In this section, we analyze the impact of generative pretraining. Since we employ Visual Autoregressive Transformer, a generative model, for depth estimation, the generative pretraining stage plays a crucial role in the model’s performance. Understanding and effectively operating in the VQ-VAE latent space is a challenging task, and the Visual Autoregressive Transformer learns to navigate this space during the pretraining phase. To evaluate the importance of pretraining, we conducted an experiment without the class-conditioned generation pretraining on ImageNet.
As shown in Table 2, the absence of pretraining leads to significantly degraded metrics across all evaluation criteria. These results demonstrate that the knowledge acquired during generative pretraining is essential for effective depth estimation. Furthermore, we hypothesize that utilizing more powerful pretraining schemes, particularly text-to-image generation on large-scale datasets like LAION-5B Schuhmann et al. (2022), could yield even better performance.
ETH TUM IBIMS
[…]
w/o ImageNet pretrain 0.392 0.46 0.561 0.522 0.338 0.301
with ImageNet pretrain 0.196 0.177 0.275 0.178 0.129 0.106
Table 2: Ablation results demonstrate that the generative pretraining of Visual Autoregressive Transformer is crucial for model performance, with metrics showing significant degradation when pretraining is omitted.
7 Discussion
@@
This work demonstrates the potential of generative autoregressive modeling for monocular depth estimation. Currently, our visual autoregressive transformer builds on pretraining, which is constrained by the the ImageNet dataset. We believe that pretraining on more extensive and diverse datasets, such as those used for text-to-image generation, could significantly enhance our model’s performance. A primary limitation of our approach is the reliance on components from the original VAR: both the VQ-VAE network and the generative pretraining of the transformer. The VQ-VAE was trained at low resolution on the relatively smaller OpenImages dataset Kuznetsova et al. (2020), in contrast to larger, more recent datasets like LAION-5B Schuhmann et al. (2022). Another potential limitation stems from the VQ-VAE training specifics: there is no guarantee that the quantization operation can accurately represent any scale that serves as a target during training. This limitation may lead to scenarios where the model converges to a local minimum due to the inherent constraints of the VQ-VAE’s learned codebook. We anticipate that upgrading to a higher-quality VQ-VAE and employing stronger pretraining could greatly benefit our method, and we identify these limitations as key directions for future research.
8 Conclusion
@@
In this paper, we tackle the depth estimation problem through an autoregressive lens, specifically adapting the visual autoregressive modeling approach Tian et al. (2024) for this task. Originally designed for class-conditioned image generation, we repurposed the Visual Autoregressive Transformer for image-conditioned depth map estimation. Our analysis highlights limitations in the standard VAR training process - teacher forcing, which leads to suboptimal accuracy on public depth benchmarks. To address these challenges, we proposed a novel training formulation, the Depth Autoregressive Refinement Task (DepthART). The Visual Autoregressive Transformer trained with DepthART showed substantial performance improvements over the teacher forcing procedure and achieved competitive or superior results on public benchmarks compared to recent methods. Our approach enhances the model’s self-refinement ability and resolves the the training-inference discrepancy issues of visual autoregressive modeling, as demonstrated through empirical evaluation.
Contribution Statement
Dmitry Senushkin - project leader